OpenStudy (anonymous):

Use properties of exponents to simplify \[\left( 27m^\frac{3}{2}n^{-6} \right)^\frac{2}{3}\]

OpenStudy (anonymous):

yeah it typed correct this time

OpenStudy (anonymous):

one at at time \[27^{\frac{2}{3}}=\sqrt[3]{27}^2=3^2=9\]

OpenStudy (anonymous):

\[(m^{\frac{3}{2}})^{\frac{2}{3}}=m^{\frac{3}{2}\times \frac{2}{3}}=m^1=m\]\]

OpenStudy (anonymous):

thye cancel each other out correct

OpenStudy (anonymous):

they even

OpenStudy (campbell_st):

use the power of a power law \[(x^a)^b= x^{a \times b}\] so its \[(3^3)^{\frac{2}{3}}\times (m^{\frac{3}{2}})^{\frac{2}{3}} \times (n^{-6})^{\frac{2}{3}}\]

OpenStudy (anonymous):

\[(n^{-6})^{\frac{2}{3}}=n^{-6\times \frac{2}{3}}=n^{-4}=\frac{1}{n^4}\]

OpenStudy (anonymous):

yes they cancel

OpenStudy (anonymous):

ok i think i already asked this question, i ment to type something different, i am so sorry, do you want me to close this and ask agian?